A 200kg object and a 500kg object are separated by 4.00m Find
(a)the net gravitational force exerted by these objects on a 50.0 kg object placed midway between them.
(b) At position (other than an infinitely remote one) can the 50.0kg object be placed so as to experience a net force of zero from the other two objects?
1
Expert's answer
2021-02-19T18:59:33-0500

(a) Let’s calculate the resultant gravitational force on a mass "m=50\ kg". According to the Newton’s law of universal gravitation, the mass "m_1=200\ kg" attracts the mass "m" with a gravitational force "F_1". The object "m_2=500\ kg" attracts the mass "m" with a gravitational force "F_2". Then, the sum of these two forces, taken algebraically (because they are along the same line) will be the net gravitational force "F_{net}":


"F_{net}=F_1+F_2,""F_{net}=-G\dfrac{m_1m}{r^2}+G\dfrac{m_2m}{r^2},""F_{net}=\dfrac{Gm}{r^2}(m_2-m_1),""F_{net}=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot50\ kg}{(2\ m)^2}\cdot(500\ kg-200\ kg)=2.5\cdot10^{-7}\ N."

(b) At a point between the two objects at a distance "d" from the "500\ kg" object, the net force on the "50\ kg" object will be zero when "F_1=F_2":


"G\dfrac{m_1m}{(r-d)^2}=G\dfrac{m_2m}{d^2},""\dfrac{m_1}{(r-d)^2}=\dfrac{m_2}{d^2},""\dfrac{d}{r-d}=\sqrt{\dfrac{m_2}{m_1}},""d=\dfrac{r\sqrt{\dfrac{m_2}{m_1}}}{1+\sqrt{\dfrac{m_2}{m_1}}},""d=\dfrac{4\ m\cdot\sqrt{\dfrac{500\ kg}{200\ kg}}}{1+\sqrt{\dfrac{500\ kg}{200\ kg}}}=2.45\ m."